1.Polling is based on the concept of random sampling. A pollster wants to know the opinion of a large population (for example, all voters in United States, which number more than 100 millions; in the 2012 election, more than 125 million people cast a ballot in the presidential election). Although it is cost-prohibitive to do a census of the entire population, he can sample a few people to get an idea of the opinion of the whole population. If his goal is to get a poll with a margin of error of about 3%, what is the minimum number of people he needs to sample?


1
Expert's answer
2021-02-03T16:35:26-0500

"n=\frac{z^2}{E^2}p(1-p)\\n=\frac{1.96^2}{0.03^2}0.5(1-0.5)=1068"